MOA-II Obsevation Log Monthly Summary 2008/05
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 13072-14418 | 1347 | 6950 | 407565 |
LMC | 2799-2868 | 70 | 386 | 102300 |
SMC | 651-675 | 25 | 50 | 11000 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 1442 | 7386 | 520865 (144h41m05s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
292 | 298 | 248 | 234 | 1074 | 222 | 230 | 443 | 1057 | 305 | 225 | 268 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
207 | 282 | 200 | 200 | 198 | 197 | 189 | 220 | 182 | 179 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
16 | 16 | 17 | 17 | 16 | 17 | 70 | 63 | 56 | 21 | 17 | 20 |
Field13 | Field14 | Field15 | Field16 |
20 | 20 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
25 | 25 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2008/05/1: Dark D3645 and D3656 are OK. Intensive observations to MOA96. / night time = 42960(sec)
2008/05/2: DARK D3659 and D3661 are OK. / night time = 43080(sec)
2008/05/3: DARK D3663 and D3664 are OK. / night time = 43200(sec)
2008/05/4: DARK D3665 and DD3666 are OK. / night time = 43320(sec)
2008/05/5: DARK D3668 and D3669 are OK. Intensive observations to MOA151. / night time = 43500(sec)
2008/05/6: DARK D3673 and D3674 are OK. / night time = 43620(sec)
2008/05/7: DARK D3676 and D3690 are OK. Anomaly (?) in MOA171. / night time = 43740(sec)
2008/05/8: No observation. / night time = 43860(sec)
2008/05/9: Intense observations of MOA171 / night time = 43980(sec)
2008/05/10: MOA144 and 171 / night time = 44100(sec)
2008/05/11: MOA171, 144 / night time = 44220(sec)
2008/05/12: / night time = 44340(sec)
2008/05/13: MOA195 and 197 / night time = 44460(sec)
2008/05/14: MOA198 and 199 / night time = 44580(sec)
2008/05/15: MOA159, 198, 199 / night time = 44700(sec)
2008/05/16: / night time = 44760(sec)
2008/05/17: / night time = 44820(sec)
2008/05/18: MOA159 Anomaly in MOA177 / night time = 44940(sec)
2008/05/19: MOA159, 177 anomaly / night time = 45060(sec)
2008/05/20: / night time = 45180(sec)
2008/05/21: No obs / night time = 45240(sec)
2008/05/22: No observation. Powe cut form 23:00 to morning / night time = 45420(sec)
2008/05/23: No observation. / night time = 45420(sec)
2008/05/24: No observation. / night time = 45480(sec)
2008/05/25: / night time = 45600(sec)
2008/05/26: Dark3759(60) Dark3760(300) has a little noise on ccd2. / night time = 45600(sec)
2008/05/27: / night time = 45720(sec)
2008/05/28: / night time = 45780(sec)
2008/05/29: / night time = 45900(sec)
2008/05/30: / night time = 45900(sec)
2008/05/31: / night time = 45900(sec)
TOTAL NIGHT TIME: 1384380(sec)
OPERATING RATE: 59.0 %
GB: 49.5 %
LMC: 8.5 %
SMC: 0.9 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month